OpenStudy (anonymous):

lim [cos(pi/2 +h) - cos(pi/2)]/h h-->0

OpenStudy (anonymous):

use the addition angle formula for cosine

OpenStudy (raden):

you can use the identity : cos(pi/2 +h) = -sin(h) and dont worry, that cos(pi/2) = 0

OpenStudy (raden):

so, the limit becomes lim h->0 -sin(h)/h = ... ?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

how can you get there using a theorm like l'hopitals or trapezoid or soemmthign along those lines?

OpenStudy (raden):

this question just basic, we neednt l'hopital rule.. just defenition that : lim h->0 sin(x)/x = 1 and lim h->0 tan(x)/x = 1

OpenStudy (anonymous):

so how do we solve it?

OpenStudy (raden):

lim h->0 -sin(h)/h = - lim h->0 sin(h)/h = - (1) = -1

OpenStudy (anonymous):

where did the sin(h) come from?

OpenStudy (anonymous):

Use the definition of the derivative to get \[ \lim_{h->0} \frac {f(a+h)-f(a)}h= f'(a)\\ \text { with }\\ a=\frac \pi 2\\ f(x)=\cos(x)\\ f'(x)=-\sin(x)\\ f'(a)=-\sin(\pi/2)=-1 \]

OpenStudy (anonymous):

The limit is -1

OpenStudy (anonymous):

I apologize but I am still confused. Can you help me plug it in because I don't understand what you are doing?

OpenStudy (anonymous):

lim [cos(pi/2 +h) - cos(pi/2)]/h h-->0